A refrigeration system of 10.5 tonnes capacity at an evaporator tem￾perature of – 12°C and a condenser temperature of 27°C is needed in a food storage locker. The refrigerant ammonia is sub-cooled by 6°C before entering the expansion valve. The vapour is 0.95 dry as it leaves the evaporator coil. The compression in the compressor is of adiabatic type. Using p-h chart find : (i) Condition of volume at outlet of the compressor (ii) Condition of vapour at entrance to evaporator (iii) C.O.P. (iv) Power required, in kW. Neglect valve throttling and clearance effect
1
Expert's answer
2021-10-31T17:35:51-0400

"_TE=T_{satPe}=\u221212\u2103,T_C=T_{satPc}=28\u2103"


Capacity=10 TR, x1=0.95


"h_1=h_{f\u221212\u2103}+x_1\u00d7h_{fg\u221212\u2103}"


"h_{fg−12℃}=h_g−h_f=1447.74−144.929

=1302.811~\frac{kJ}

{kg}"


(from psychrometric chart booklet for N3)


"h_1=144.929+0.95×1302.811

=1382.6~\frac{kJ}

{kg}"


Condition of refrigerant at outlet of compressor


"T_2=67°C,h_2=1598~\frac{kJ}{kg},v_2=0.15~\frac{m^3}{kg},p_2=10.5~ bar"

Condition of refrigerant at inlet to evaporator


"T_4=−12°C,h_4=310~\frac{kJ}{kg},p_4=2.65 ~bar"

COP of given system = "\frac{R.E.}{W_c}=\frac{h_1−h_4}{h_2−h_1}=4.98"

Also, 

"Q\u0307_a=m\u0307_R\u00d7(h_1\u2212h_4)"

For 10 TR capacity, "10x3.5=m\u0307_R\u00d7(1382.6\u2212310)"

mass flow rate, "ṁ_R=0.0326~\frac{kg}{sec}"

Now, power required = "m\u0307_R\u00d7(h_2\u2212h_1)=7.0287 ~kW."